OpenStudy (anonymous):

Variable y varies directly with x^2, and y = 196 when x = 7. Enter the constant of variation in the quadratic variation equation. y = _________ x^2

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

or anybody else Remember I always give FAN and MEDAL

OpenStudy (danjs):

hi

OpenStudy (danjs):

y = k*x^2

OpenStudy (danjs):

(x,y) = (7,196)

OpenStudy (danjs):

196 = k*(7)^2

OpenStudy (danjs):

k = 196/49

OpenStudy (danjs):

\[y = 4 * x^2\]

OpenStudy (danjs):

Just remember, varies directly means y = k*x and indirectly proportional is, \[y=k*\frac{ 1 }{ x}\]